OpenStudy (anonymous):

Hanna makes her own furniture polish using twenty-five percent vinegar and seventy-five percent olive oil. Write and solve an equation to determine how much total solution she made if she used 2 cups of vinegar. Show your work.

OpenStudy (anonymous):

What you want to do is find out what 2 is 25% of. To do this you can make 25% into a fraction, 1/4, and now we know that 2 = 1/4. Knowing this we can multiply both sides by 4, so now, 8 = 1, where 8 represents how many cups of liquid are in 1 mixture. Now we know that there are a total of 8 cups of liquid in this mixture and 2 of the 8 cubs are vinegar. You can multiply 8 by 75%, or .75, to find the amount of oil put in. 8*.75=6 So there are 8 cups of total solution :3

OpenStudy (anonymous):

You can make sure by adding the total oil and vinegar 6+2=8 yay

OpenStudy (anonymous):

thanks